OpenStudy (anonymous):

Using the domain values {-1, 0, 4}, find the range values for relation y=2x -7

OpenStudy (precal):

domain is the x values just sub those in, one at a time y=2(-1)-7 y=-2-7 y=-9 that is your first one do the rest in the same fashion